HUBBLE SPACE TELESCOPE DAILY REPORT #1160 PERIOD COVERED: 0700 EST Daily Status Report as of 166/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1 5438 (Io's Atmosphere: UV Spectroscopy of Neutral and Molecular Components) The High Resolution Spectrograph (Side One) was used to make far ultraviolet spectrographic observations of Jupiter's moon Io. The observations were executed as scheduled, and the data remains on Science Tape Recorder #3. The data will be analyzed later. No anomalies were noted. 1.2 Completed GHRS/1 5541 (GHRS Side 1 Dark Noise Monitor: Cycle 4) The High Resolution Spectrograph (Side One) was used to make routine detector dark noise calibration observations. These were internal tests and did not require any special pointing of the telescope. The observations were executed as scheduled, and the data remains on Science Tape Recorder #3. No anomalies were noted. 1.3 Completed FOS/RD 5536 (Cycle 4 FOS Wavelength Calibration: Internal/External Offsets) The Faint Object Spectrograph (Red Side) made both internal and external wavelength calibration observations. HD207757 was the external source. The observations were executed as scheduled, and the data remains on Science Tape Recorder #3. No anomalies were noted. 1.4 Completed Three Sets of WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were executed as scheduled, and the data remains on Science Tape Recorder #3. The data will be analyzed later. No anomalies were noted. 1.5 Completed WF/PC-2 5629 (WF/PC-2 Cycle 2 Photometric Calibration Monitor: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and the data remains on Science tape Recorder #3. No anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 5465 (WF/PC-2 Imaging of Two Galaxies at Z=1.2) The WF/PC-2 was used to make deep images of galaxy cluster 3C324 at Z=1.2. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earthflats, Large Set) The WF/PC-2 was used to make a set of Earth flat calibration images using different filters in order to construct high quality flat fields of the OTA illumination pattern. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled :10 Successful :10 2.3 Significant HSTARs: None 2.4 Operations Notes: Another uneventful operational day was experienced with all HST subsystems reporting anomaly-free activities. 3.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-